Pierre paid $126 for a coat that was listed for $180. What rate of discount did he receive?
Found 2 solutions by fractalier, JBarnum:
Answer by fractalier(6550)

Rate of discount = the amount of discount / original price =
= (180 - 126) / 180 = 54 / 180 = .3 = 30%

Answer by JBarnum(2146)

180-126=54
54/180=0.3
0.3*100=30% discount